OpenStudy (anonymous):

The hypotenuse of a right triangle is 5 inches long. The shorter leg is 1 inch shorter than the longer leg. Find the lengths of the triangle.

OpenStudy (ipwnbunnies):

You know the Pythagorean Theorem, right? \[a^2 + b^2 = c^2\] Well, let 'x' be the length of the longer leg. \[x^2 + (x-1)^{2} = 5^2\]